### v2

# Optical Injection and Detection of Long-Lived Interlayer Excitons in van der Waals Heterostructures

Abstract
Interlayer excitons in semiconducting bilayers separated by insulating hBN layers constitute a promising platform for investigation of strongly correlated bosonic phases. Here, we report an optical method for the generation and characterization of long-lived interlayer excitons. We confirm the presence of tightly bound interlayer excitons by measuring 1s and 2s intralayer excitons in each layer concurrently. Using a pump-probe technique, we find interlayer exciton lifetimes up to , increasing with the thickness of the hBN. With optical access to long-lived interlayer excitons, our approach provides a new route to explore degenerate Bose–Fermi mixtures of excitons and itinerant electrons with high spatial and temporal resolution.
Transition metal dichalcogenide (TMD) heterobilayers have recently emerged as a promising platform for creating long-lived interlayer excitons, bound electron-hole pairs, whose constituents reside in different layers. With their large binding energy and a built‑in out‑of‑plane dipole, they offer a route to investigate many-body phenomena ranging from superfluidity of dipolar excitons [1, 2, 3, 4] to Bose–Fermi mixtures exhibiting exotic electron pairing mechanisms [5, 6] in van der Waals heterostructures. However, accessing these collective states depends on controlling the lifetime, density, and interactions of interlayer excitons [7, 3].
Early demonstrations of long-lived interlayer excitons relied on the electrical injection of charge carriers into separate TMD layers [8, 9, 10, 11]. Capacitance spectroscopy and Coulomb‑drag measurements provided compelling evidence for interlayer exciton formation. However, these measurements offer no information on time dynamics and spatial distributions. In parallel, optical pumping in aligned MoSe2–WSe2 bilayers, either in direct contact or separated by one or two hexagonal boron nitride (hBN) layers [12, 13, 14, 15] has been explored. Photoluminescence (PL) from such devices has revealed lifetimes up to [16]. However, achieving longer lifetimes and higher densities requires thicker hBN spacer layers to suppress residual interlayer tunneling and radiative recombination [17]. This, in turn, significantly reduces PL emission, rendering direct optical measurements of the interlayer exciton dynamics untenable.
Here, we demonstrate an optical scheme that overcomes the limitations of both electrical and PL-based injection and detection schemes. While inserting hBN spacer with up to seven layers to suppress interlayer tunneling, we inject high densities of electrons and holes into separate MoSe2 and WSe2 layers via nonresonant optical pumping, without requiring layer-selective contacts or complex gating geometries [8, 9]. Using time-resolved reflection spectroscopy and the 2s exciton resonance as a spectroscopic probe, we observe interlayer exciton formation and relaxation dynamics in samples with negligible interlayer exciton PL. Under these conditions, we achieve interlayer exciton densities up to and lifetimes of . This combination of high density, long lifetime, and optical control opens the door to studying strongly interacting interlayer exciton gases and exploring collective excitonic phenomena.
We investigate three TMD bilayer devices with varying hBN spacers (see Fig. 1a). In the main text, we present data from Devices 1 and 2, both MoSe2–WSe2 bilayers with hBN spacers of 1- and 3–5-layers, while results from Device 3 (MoS2–WSe2, 5–7 layer hBN spacer), are provided in the Supplemental Information (SI). In contrast to previous optical studies, all TMD layers within a device are angle-misaligned, suppressing interlayer hybridization, and resulting in negligible interlayer PL 111Device 1 shows very weak interlayer PL, whereas no interlayer PL is observed in Devices 2 and 3; see Fig. S2. To determine the charge configuration in the individual layers, we perform reflection spectroscopy. The upper panel of Fig. 1b shows spectra of the MoSe2 and WSe2 1s exciton resonances as a function of the gate voltage (controls the chemical doping of the layers, see SI for details), measured in Device 1. Upon injection of itinerant charge carriers, the 1s exciton evolves into two distinct resonances (Fig. 1b): the repulsive polaron (RP) and the attractive polaron (AP) [19]. The left (right) side of the blue (red) dashed line in Fig. 1b corresponds to hole (electron) injection into the WSe2 (MoSe2) layer. The voltage range between the dashed lines marks the charge-neutral regime, where neither layer is doped.
Unlike previous experiments that bridged this energy gap using large electric fields and an interlayer bias [8, 9, 11, 10], we employ nonresonant optical pumping to inject charge carriers into the TMD bilayer. We illuminate the sample with a continuous-wave laser at 635 nm, above the bandgap (Fig. 1c) with pump and probe focused onto the same spot. Remarkably, as we show below, this optical doping technique enables simultaneous electron- and hole-doping in opposite layers even for devices with thicker hBN spacers, overcoming the intrinsic band-offset limitations without the need for applying an electrical bias. Fig. 1d displays the reflection spectra under continuous pumping at laser power . Here, the charge-neutral region vanishes, and we observe simultaneous electron doping in the MoSe2 layer and hole doping in the WSe2 layer. This is indicated by the reversal in the positions of the red and blue dashed lines 222We emphasize that the two types of doping can be spectrally distinguished in WSe2 [35]: the presence of a single attractive polaron (AP) resonance in the WSe2 spectrum unequivocally indicates hole doping, since electron doping of WSe2 leads to two AP features associated with the singlet and triplet trions. In contrast, MoSe2 displays a single AP peak under both electron and hole doping. Additional confirmation comes from Device 3, a MoS2/WSe2 heterostructure, where the MoS2 layer clearly exhibits two AP features under electron doping, consistent with our interpretation of the spectra in Devices 1 and 2 (see Fig. S6)..
To visualize the overall charge configuration, Fig. 1e shows the sum of the peak 1s X (RP) reflection contrast of MoSe2 and WSe2 layers normalized to the individual reflection contrast of each layer in the absence of charges () as a function of and (see SI for details). The yellow region labeled as (i,i) and outlined by a black-dotted contour denotes the approximate charge-neutral regime, where both excitons reach maximum contrast. At higher , a dark blue region appears above a light gray dotted contour line, indicating dual-layer doping with opposite charge carriers (e,h). In the surrounding regions, either WSe2 is hole doped (i,h) or MoSe2 is electron doped (e,i), with the opposite layer remaining charge neutral. Details on the asymmetry between hole and electron doping are provided in the SI.
Next, we investigate whether the injected electrons and holes form bound interlayer exciton states, or remain unbound. Prior theoretical work [21, 9] demonstrated that the modification of the 1s exciton-polaron resonances can be used to differentiate between bound and unbound electron-hole pairs; however, the associated spectral shifts are rather small and can be masked by exciton line broadening. The origin of this relatively weak dependence is the small Bohr radius of 1s intralayer excitons, rendering them insensitive to the charges in the neighboring TMD layer (see the sketch in the upper panel of Fig. 2a). In contrast, Rydberg exciton states, such as the 2s state, exhibit a significantly larger Bohr radius (lower panel of Fig. 2a), making them more sensitive to the surrounding electrostatic environment [22]. As such, the 2s exciton reflection contrast can be expected to respond differently depending on whether the charge carriers are (tightly) bound as interlayer excitons or remain free unbound particles.
Spectrally, we can resolve MoSe2 and WSe2 2s and 3s excitons in the charge-neutral regime (Fig. 2b). In the case of free carriers residing in the same layer, 2s excitons exhibit a blue shift and loss of oscillator strength even for small carrier densities. When charges are located in the opposite layer, the 2s-resonance still quickly disappears, but part of its oscillator strength is transferred to a new red-shifted resonance which arises from the hybridization with the 2p-states [22, 23, 24, 25].
Interestingly, we observe that in the regime where both layers are doped under finite , the MoSe2 and WSe2 2s exciton features persist with reduced intensity and only a small spectral redshift (see SI). To quantify this behavior across the charge map, we extract the WSe2 2s contrast, defined as the peak-to-peak amplitude of the dispersive Lorentzian lineshape associated with the WSe2 2s exciton (Fig. 2c).
Figures 2d–f show differentiated spectral linecuts taken at different positions in the phase diagram. In the (e,h) regime (black diamond), the 2s exciton contrast remains finite, even though the charge densities match those in the red and blue diamond regions, where the 2s resonance vanishes. Moreover, the 2s resonance energy in the (e,h) region coincides with that at charge neutrality (green diamond). This robustness of the 2s resonance in the presence of dual-layer doping stands in stark contrast to its drastic modification due to free charges in either layer, and points to a distinct phase: the formation of interlayer excitons. Because interlayer excitons are charge neutral and do not generate strong in-plane electric fields, they do not directly hybridize the 2s and 2p-states. Instead, the leading-order interaction between the 2s and interlayer excitons is a weak attractive van der Waals interaction (see SI), leading to a small density-dependent redshift.
We note that the persistence of the 2s was consistently observed at different spots within a device as well as across multiple devices, indicating that spatial inhomogeneities are unlikely to account for the observed feature. Since the measured electron and hole trion binding energies in the two layers agree well with previously studied monolayer devices, we can also rule out strong trapping of the optically introduced charges [26] as an explanation for the robustness of 2s excitons.
We further observe that the 2s exciton contrast diminishes progressively with increasing . This behavior is expected: as the interlayer exciton spacing decreases and approaches the 2s Bohr radius, the 2s state will cease to be bound. The fact that the 3s exciton resonance vanishes at even lower interlayer exciton densities supports our explanation (Fig. S7). We emphasize that the 2s signal quenches at densities far below the expected interlayer exciton Mott density of [21, 8].
Figures 3a and 3b show maps of the electron and hole densities, and , as functions of and (see SI). From these data we extract the minimum of () and the excess charge density , plotted in Figures 3c and 3d.
In the region where the 2s contrast is visible and the excess charge approaches zero (pink dashed contour), all carriers are paired and thus form interlayer excitons. Thus, the interlayer exciton density is equal to . Outside the pink contour, interlayer excitons may still be present alongside free carriers, but the 2s spectroscopic signature lacks sufficient contrast to confirm their presence. Here, provides an upper bound on the attainable interlayer exciton density. In Device 1, where the MoSe2 and WSe2 layers are separated by a monolayer of hBN, the min reaches up to . By contrast, in Device 2, which features an hBN spacer thickness of 3–5 layers, min reaches (Fig. S10).
Our optical pumping scheme provides access to interlayer exciton dynamics. We probe these dynamics using time-resolved reflection spectroscopy, with synchronized modulation of the pump and probe beams that allows for programmable delays between pulses (see SI for details). In Fig. 4, we present the evolution of on a logarithmic scale as a function of pump–probe delay, enabling us to quantify the lifetime of the interlayer exciton state at a representative point in the phase diagram, where is finite, the total excess charge is negligible, and the WSe2 2s exciton is present. The decay of the signal with increasing time delay reflects the relaxation of interlayer excitons after the pump laser is turned off. Although the decay is clearly nonexponential, we extract a 1/e lifetime of for Device 1, which contains a monolayer hBN spacer. In contrast, a significantly longer lifetime of is observed for Device 2, which has a thicker hBN spacer of 3–5 layers.
This prolonged lifetime is consistent with expectations: increased spacer thickness suppresses interlayer tunneling and reduces the radiative decay rate, thereby extending the interlayer exciton lifetime [16]. Surprisingly, the decay becomes faster at lower carrier densities 333The parabolic trend in the logarithmic plot in Fig. 4 is clearly visible; similar nonexponential time dependence has also been observed by Cutshall et al. [16] in a time-resolved PL measurement. We speculate that the shortened lifetime may stem from interlayer excitons becoming more tightly bound as their density is lowered: Since the radiative decay rate scales inversely with the square of the Bohr radius, we would expect the lifetime to become shorter. Lastly, we emphasize that our measurements cannot rule out a dominant, sample-dependent, nonradiative decay mechanism.
The generation of interlayer excitons within a focused optical spot, combined with the ability to probe their dynamics on picosecond timescales [28], provides a powerful platform for exploring degenerate dipolar excitons with high spatial and temporal resolution. Crucially, it allows spatially controlled exciton generation without the need for external electric fields, which are typically required in schemes requiring bias voltage for interlayer exciton generation [8, 9]. This capability opens the door to creating multiple excitonic reservoirs within a single device, paving the way for Josephson-like experiments [29, 30, 31]. In particular, optically inducing high-density exciton populations in two gate-defined traps [32, 33] connected by a narrow channel could provide a direct route to optically probing [34] coherent tunneling and phase dynamics between condensate regions.
Acknowledgements.
We thank Xiaobo Lu for fabricating Device 1. We thank Ivan Amelio, Haydn S. Adlong and Igor Khanonkin for inspiring discussions. This work was supported by the Swiss National Science Foundation (SNSF) under Grant No. 200021-204076. A.M.S. and A.C. acknowledge funding from an ETH Postdoctoral Fellowship. K.W. and T.T. acknowledge support from the JSPS KAKENHI (Grant Numbers 21H05233 and 23H02052), the CREST (JPMJCR24A5), JST and World Premier International Research Center Initiative (WPI), MEXT, Japan.References
- Lu et al. [2011] M. Lu, N. Q. Burdick, S. H. Youn, and B. L. Lev, Phys. Rev. Lett. 107, 190401 (2011).
- Aikawa et al. [2012] K. Aikawa, A. Frisch, M. Mark, S. Baier, A. Rietzler, R. Grimm, and F. Ferlaino, Phys. Rev. Lett. 108, 210401 (2012).
- Alloing et al. [2014] M. Alloing, M. Beian, M. Lewenstein, D. Fuster, Y. González, L. González, R. Combescot, M. Combescot, and F. Dubin, EPL 107, 10012 (2014).
- Combescot et al. [2017] M. Combescot, R. Combescot, and F. Dubin, Rep. Prog. Phys. 80, 066501 (2017).
- Zerba et al. [2024] C. Zerba, C. Kuhlenkamp, A. Imamoğlu, and M. Knap, Phys. Rev. Lett. 133, 056902 (2024).
- von Milczewski et al. [2024] J. von Milczewski, X. Chen, A. Imamoglu, and R. Schmidt, Phys. Rev. Lett. 133, 226903 (2024).
- High et al. [2012] A. A. High, J. R. Leonard, A. T. Hammack, M. M. Fogler, L. V. Butov, A. V. Kavokin, K. L. Campman, and A. C. Gossard, Nature 483, 584 (2012).
- Ma et al. [2021] L. Ma, P. X. Nguyen, Z. Wang, Y. Zeng, K. Watanabe, T. Taniguchi, A. H. MacDonald, K. F. Mak, and J. Shan, Nature 598, 585 (2021).
- Qi et al. [2023] R. Qi, A. Y. Joe, Z. Zhang, Y. Zeng, T. Zheng, Q. Feng, J. Xie, E. Regan, Z. Lu, T. Taniguchi, K. Watanabe, S. Tongay, M. F. Crommie, A. H. MacDonald, and F. Wang, Nature Communications 14, 8264 (2023).
- Nguyen et al. [2025] P. X. Nguyen, L. Ma, R. Chaturvedi, K. Watanabe, T. Taniguchi, J. Shan, and K. F. Mak, Science 388, 274 (2025).
- Qi et al. [2025] R. Qi, A. Y. Joe, Z. Zhang, J. Xie, Q. Feng, Z. Lu, Z. Wang, T. Taniguchi, K. Watanabe, S. Tongay, and F. Wang, Science 388, 278 (2025).
- Rivera et al. [2015] P. Rivera, J. R. Schaibley, A. M. Jones, J. S. Ross, S. Wu, G. Aivazian, P. Klement, K. Seyler, G. Clark, N. J. Ghimire, et al., Nat. Commun. 6, 6242 (2015).
- Jauregui et al. [2019] L. A. Jauregui, A. Y. Joe, K. Pistunova, D. S. Wild, A. A. High, Y. Zhou, G. Scuri, K. De Greve, A. Sushko, C.-H. Yu, et al., Science 366, 870 (2019).
- Shimazaki et al. [2020] Y. Shimazaki, I. Schwartz, K. Watanabe, T. Taniguchi, M. Kroner, and A. Imamoğlu, Nature 580, 472 (2020).
- Mahdikhanysarvejahany et al. [2022] F. Mahdikhanysarvejahany, D. N. Shanks, M. Klein, Q. Wang, M. R. Koehler, D. G. Mandrus, T. Taniguchi, K. Watanabe, O. L. Monti, B. J. LeRoy, et al., Nat. Commun. 13, 5354 (2022).
- Cutshall et al. [2025] J. Cutshall, F. Mahdikhany, A. Roche, D. N. Shanks, M. R. Koehler, D. G. Mandrus, T. Taniguchi, K. Watanabe, Q. Zhu, B. J. LeRoy, et al., Sci. Adv. 11, eadr1772 (2025).
- Snoke [2011] D. Snoke, Adv. Condens. Matter Phys. 2011, 938609 (2011).
- Note [1] Device 1 shows very weak interlayer PL, whereas no interlayer PL is observed in Devices 2 and 3; see Fig. S2.
- Sidler et al. [2017] M. Sidler, P. Back, O. Cotlet, A. Srivastava, T. Fink, M. Kroner, E. Demler, and A. Imamoglu, Nat. Phys. 13, 255 (2017).
- Note [2] We emphasize that the two types of doping can be spectrally distinguished in WSe2 [35]: the presence of a single attractive polaron (AP) resonance in the WSe2 spectrum unequivocally indicates hole doping, since electron doping of WSe2 leads to two AP features associated with the singlet and triplet trions. In contrast, MoSe2 displays a single AP peak under both electron and hole doping. Additional confirmation comes from Device 3, a MoS2/WSe2 heterostructure, where the MoS2 layer clearly exhibits two AP features under electron doping, consistent with our interpretation of the spectra in Devices 1 and 2 (see Fig. S6).
- [21] I. Amelio, N. D. Drummond, E. Demler, R. Schmidt, and A. Imamoglu, Phys. Rev. B 107, 155303, publisher: American Physical Society.
- Xu et al. [2020] Y. Xu, S. Liu, D. A. Rhodes, K. Watanabe, T. Taniguchi, J. Hone, V. Elser, K. F. Mak, and J. Shan, Nature 587, 214 (2020).
- Popert et al. [2022] A. Popert, Y. Shimazaki, M. Kroner, K. Watanabe, T. Taniguchi, A. Imamoğlu, and T. Smoleński, Nano Letters 22, 7363 (2022).
- Kim et al. [2025] J. Kim, H. Dery, and D. Van Tuan, Phys. Rev. B 112, L041301 (2025).
- [25] A. Christianen, A. M. Seiler, A. Tüğen, and A. İmamoğlu, in preparation .
- Kiper et al. [2025] N. Kiper, H. S. Adlong, A. Christianen, M. Kroner, K. Watanabe, T. Taniguchi, and A. İmamoğlu, Phys. Rev. X 15, 011049 (2025).
- Note [3] The parabolic trend in the logarithmic plot in Fig. 4 is clearly visible; similar nonexponential time dependence has also been observed by Cutshall et al. [16] in a time-resolved PL measurement.
- Uto et al. [2024] T. Uto, B. Evrard, K. Watanabe, T. Taniguchi, M. Kroner, and A. İmamoğlu, Phys. Rev. Lett. 132, 056901 (2024).
- Khorana [1969] B. Khorana, Phys. Rev. 185, 299 (1969).
- Lagoudakis et al. [2010] K. G. Lagoudakis, B. Pietka, M. Wouters, R. André, and B. Deveaud-Plédran, Phys. Rev. Lett. 105, 120403 (2010).
- Abbarchi et al. [2013] M. Abbarchi, A. Amo, V. Sala, D. Solnyshkov, H. Flayac, L. Ferrier, I. Sagnes, E. Galopin, A. Lemaître, G. Malpuech, et al., Nat. Phys. 9, 275 (2013).
- Rapaport et al. [2005] R. Rapaport, G. Chen, S. Simon, O. Mitrofanov, L. Pfeiffer, and P. Platzman, Phys. Rev. B 72, 075428 (2005).
- Gärtner et al. [2007] A. Gärtner, L. Prechtel, D. Schuh, A. Holleitner, and J. Kotthaus, Phys. Rev. B 76, 085304 (2007).
- Regan et al. [2020] E. C. Regan, D. Wang, C. Jin, M. I. Bakti Utama, B. Gao, X. Wei, S. Zhao, W. Zhao, Z. Zhang, K. Yumigeta, et al., Nature 579, 359 (2020).
- Wang et al. [2020] T. Wang, Z. Li, Z. Lu, Y. Li, S. Miao, Z. Lian, Y. Meng, M. Blei, T. Taniguchi, K. Watanabe, et al., Phys. Rev. X 10, 021024 (2020).